# Assessing the Human Likeness of AI-Generated Counterspeech

### v1

@@ Abstract
Counterspeech is a targeted response to counteract and challenge abusive or hateful content. It can effectively curb the spread of hatred and foster constructive online communication. Previous studies have proposed different strategies for automatically generated counterspeech. Evaluations, however, focus on the relevance, surface form, and other shallow linguistic characteristics. In this paper, we investigate the human likeness of AI-generated counterspeech, a critical factor influencing effectiveness. We implement and evaluate several LLM-based generation strategies, and discover that AI-generated and human-written counterspeech can be easily distinguished by both simple classifiers and humans. Further, we reveal differences in linguistic characteristics, politeness and specificity.
@@
Xiaoying Song1, Sujana Mamidisetty1, Eduardo Blanco2, Lingzi Hong1 1College of Information, University of North Texas 2Department of Computer Science, University of Arizona
Triggering warning: Read with caution. Examples in this paper may present toxic languages.
@@ 1 Introduction
Table 1: A hate speech post from Reddit and AI-generated counterspeech reply. In this paper, we compare AI-generated and human-written counterspeech replies and conclude that are fundamentally different.
@@
Research has been devoted to curating counterspeech Qian et al. (2019); Chung et al. (2021); Yu et al. (2022). Human-written counterspeech can positively affect bystanders and encourage more civil online interaction  Zhao et al. (2024). However, human involvement is often time-consuming and costly. Generative AI models have also been developed to automatically generate counterspeech Zhu and Bhat (2021); Tekiroglu et al. (2022); Saha et al. (2022); Hassan and Alikhani (2023); Hong et al. (2024). Automatic generation has the potential to mimic human-written counterspeech and assist in combating hate Bail (2024).
@@
However, generative AI models may struggle to accurately understand semantic nuances, leading to misunderstandings and even backfiring Luger and Sellen (2016); Hadi (2019); Toader et al. (2019). While these systems attempt to convey empathy like humans, sometimes they do not appropriately match the user’s emotional state or the conversation context  Han et al. (2022); Chung et al. (2021). Table 1 shows an example of counterspeech generated by prompt with Llama2.
While it aims to mimic human responses, the generated counterspeech may lack the depth of empathy and cultural sensitivity crucial for effective communication Benesch et al. (2016); Hangartner et al. (2021).
@@
Human likeness is an important factor in crafting counterspeech. Studies find that more human-like counterspeech tends to be more effective Pinochet et al. (2024). The tone and form of counterspeech can significantly affect the effectiveness in influencing behaviors of users Buerger (2021). Generated counterspeech that embodies human characteristics including empathy, humor, and respect can foster emotional attachments among users Glikson and Woolley (2020), potentially making counterspeech more acceptable and reliable Go and Sundar (2019a); Jiang et al. (2023b). Human-like responses are more likely to express understanding and social connection Go and Sundar (2019a), contribute to fostering genuine interactions, and build trust Fenwick and Molnar (2022); Cheng and Pan (2021).
@@
On the other hand, there are ethical concerns if AI-generated and human-written counterspeech are indistinguishable Hua et al. (2024). AI models may be used for manipulative purposes such as swaying opinions not aligned with specific agendas Karnouskos (2020). AI models may also generate biased responses that could unintentionally reinforce stereotypes or cause harm Lucy and Bamman (2021); Kenthapadi et al. (2023). Evaluating the human likeness of AI-generated counterspeech is important because it brings awareness to these issues and could help mitigate them.
@@
Previous studies have focused on evaluating the relevance, quality, toxicity, persuasiveness, and effectiveness of counterspeech Saha et al. (2022); Jiang et al. (2023a); Hong et al. (2024). They have also studied linguistic aspects using metrics such as BLEU, METEOR, BERTScore, GRUENV scores, type-token and distinct-n Chung et al. (2019); Qian et al. (2019); Chung et al. (2021); Zhu and Bhat (2021); Tekiroglu et al. (2022); Jiang et al. (2023a). However, few have conducted thorough evaluations regarding human likeness.
@@
We implement several state-of-the-art counterspeech generation strategies with LLMs, including prompting for one counterpeech reply(Prompt), prompt for multiple replies and select the best(Prompt and Select), fine-tune LLMs with existing counterspeech datasets(Fine-tune), and outcome-constrained LLMs.(Constrained). We also collect human-written counterspeech, including both examples from actual user-generated content (i.e., genuine content from Reddit) and crowd workers (i.e., counterspeech generated on demand).
Armed with AI-generated and human-written counterspeech, we build authorship attribution models to identify which AI models generate more human-like counterspeech. The higher the classification accuracy, the less human-like the AI-generated counterspeech is. We further deploy a human annotation task to validate and check whether humans can discern the differences between AI-generated and human-written counterspeech. Additionally, we analyze linguistic characteristics, politeness, and specificity of the counterspeech generated by various methods to understand the differences amongst counterspeech depending on the source (types of AI-generated and human-written).
Our study shows that AI-generated and human-written counterspeech can be easily distinguished by both classifiers and humans. Additionally, AI-generated counterspeech shows (a) significant differences in linguistic characteristics, and (b) is more polite and less specific.
@@
Domain Size Context Generation Model Training Methods
@@ 2 Related Work
Counterspeech Generation Several studies employ crowd workers or experts to curate counterspeech Chung et al. (2019); Qian et al. (2019); Yu et al. (2022). These human-written datasets have been widely utilized for training counterspeech models in subsequent research Tekiroglu et al. (2022); Saha et al. (2022); Gupta et al. (2023). Hybrid approaches that combine human annotations and generative models have been proposed, resulting in counterspeech corpora such as CONAN Chung et al. (2019) and multiCONAN Fanton et al. (2021). Advanced generative models have been developed for more sophisticated generation of counterspeech, for example, the prompt and select method has been shown to generate more diverse and relevant counterspeech Zhu and Bhat (2021). Other approaches include constraining models for generating polite Saha et al. (2022), intent-based Gupta et al. (2023), or outcome-oriented counterspeech Hong et al. (2024). This study (a) implements several representative models to generate counterspeech and (b) conducts an extensive evaluation focusing on human likeness.
@@
Most curated datasets consist of texts without additional context El-Sayed and Nasr (2023); Buz et al. (2024); An et al. (2023). Few include context and involve sequences or groups of related texts Ji et al. (2024); Prajapati et al. (2024). To distinguish between AI-generated and human-written texts, many researchers conduct linguistic analyses and authorship attribution, supplemented with human evaluations Zhou et al. (2023); Ma et al. (2023); Buz et al. (2024); Ji et al. (2024); Prajapati et al. (2024). Our study is among the few to work in a dialogue setting, where texts are dialogue turns countering a hate speech—AI-generated following four strategies or human-written from two sources.
3 Counterspeech Data Curation
@@ 3.1 Human-written Counterspeech
Human-written counterspeech consists of replies to hateful content written by humans. We consider counterspeech written by both genuine Reddit users (i.e., users that post out of their own will) and crowd workers who are tasked with writing counterspeech. Genuine human-written counterspeech is collected from Reddit. First, we use keyword- and community-based sampling methods to retrieve hate speech posts from 42 subreddits (Appendix A) with a higher prevalence of hate via the Pushshift API. This step results in 27,491 hate speech posts and their replies. Second, we identify replies that are counterspeech with three BERT classifiers individually fine-tuned with three existing counterspeech corpora Qian et al. (2019); Chung et al. (2021); Yu et al. (2022). We consider a reply to be counterspeech if the three classifiers indicate so. This step results in 14,973 (hate speech, counterspeech) pairs from Reddit.
Counterspeech written by crow workers is collected from the Benchmark dataset Qian et al. (2019). This corpus includes hate speech posts paired with counterspeech replies written by crowd workers on demand. We identify 14,208 valid (hate speech, counterspeech) pairs in Benchmark.
@@ 4 Evaluation Methods
We conduct evaluations comparing human-written and AI-generated counterspeech accounting for four major categories: human likeness, linguistic differences, politeness and specificity.
@@
If it is easy to differentiate between human-written and AI-generated counterspeech, we can conclude that the latter is not human-like. To this end, we develop BERT classifiers with our curated dataset (Section 3). The dataset is divided into an 80/20 split for training and testing. Specifically, we create five binary classifiers (human-written or AI-generated) depending on what is included in the AI-generated counterspeech: all or only the counterspeech generated by one of the four strategies (see Appendix B for details).
In addition, we conduct a human validation for deeper insights. We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask human annotators to label whether the counterspeech has human-written or AI-generated. Two research assistants complete the annotation process. Table 3 presents the agreement rate and Cohen’s Kappa score. Both indicate high reliability, meaning that the task is straightforward for humans.
@@
The level of politeness assesses the degree of respectfulness and courtesy. It provides an complementary perspective to understand the differences between AI-generated and human-written counterspeech. We build a politeness prediction model, a BERT model fine-tuned with the dataset by Saha et al. (2022). The model achieves an F1 score of 0.91. We use this model to predict the politeness level of the counterspeech replies on a scale of 0 to 7. A higher score indicates more politeness.
@@ 5.4 Specificity
The counterspeech generated by Prompt and Prompt and Select models exhibits a similar distribution with a moderate specificity level of 3 (Figure 2(b)). The Constrained model-generated counterspeech has a median specificity score of 1.5, indicating that it lacks specificity and provides a more general reply. Comparatively, the counterspeech by Fine-tune models shows high variances in the specificity scores, but most scores are between 3 and 4, indicating counterspeech more targeted to the hate speech post at hand.
@@ 7 Conclusion
We fully consider the potential risks and benefits of our study. First, we collect data from Reddit, a public forum that makes data available to third parties. We have masked users’ names and identities before analysis. Second, our study employs human annotators to evaluate counterspeech, their names and identities are encrypted to avoid the identification of annotators. Third, we provide various AI-based methods to generate counterspeech that embody human-likeness. There is a concern that some generated responses are indistinguishable from those created by humans, raising ethical considerations. We acknowledge the potential risks these methods may cause. However, the benefits will outweigh such risks, for example, using these generations in adversarial learning to develop more robust models for identifying AI-generated content.
@@ Acknowledgements
We acknowledge the use of code-writing assistance GitHub Copilot. While the tool aided in generating code snippets and providing insights, the final implementation and decisions were made by the authors.
@@ Appendix C Annoatation Details
0-Very Impolite Counterspeech is percieved as rude or unfriendly The study was b**.

### v2

@@ Abstract
Counterspeech is a targeted response to counteract and challenge abusive or hateful content. It effectively curbs the spread of hatred and fosters constructive online communication. Previous studies have proposed different strategies for automatically generated counterspeech. Evaluations, however, focus on relevance, surface form, and other shallow linguistic characteristics. This paper investigates the human likeness of AI-generated counterspeech, a critical factor influencing effectiveness. We implement and evaluate several LLM-based generation strategies, and discover that AI-generated and human-written counterspeech can be easily distinguished by both simple classifiers and humans. Further, we reveal differences in linguistic characteristics, politeness, and specificity. The dataset used in this study is publicly available for further research  [^1] https://github.com/oliveeeee25/counterspeech_eval_humanlike.
@@
Xiaoying Song1 Sujana Mamidisetty2 Eduardo Blanco3 Lingzi Hong1 1College of Information, University of North Texas 2TAMS, University of North Texas 3Department of Computer Science, University of Arizona {xiaoyingsong, sujanamamidisetty}@my.unt.edu eduardoblanco@arizona.edu lingzi.hong@unt.edu
Trigger warning: Read with caution. Examples in this paper may present toxic languages.
@@ 1 Introduction
Table 1: A hate speech post from Reddit and counterspeech generated by prompt with Llama2. We compare AI-generated and human-written counterspeech and conclude they are fundamentally different.
@@
Recent studies have explored methods for generating counterspeech Qian et al. (2019); Chung et al. (2021). Human-written counterspeech is often contextually appropriate and can positively affect bystanders and encourage civil online interactions  Zhao et al. (2024). However, human involvement is time-consuming and costly. Models have been developed to automatically generate counterspeech Zhu and Bhat (2021); Tekiroglu et al. (2022); Hong et al. (2024). The provision has the potential to mimic human-written counterspeech and assist in combating hate Bail (2024).
@@
However, counterspeech generation models may struggle to accurately understand semantic nuances, leading to misunderstandings and even backfiring Luger and Sellen (2016); Hadi (2019); Toader et al. (2019). Sometimes they do not appropriately match the user’s emotional state or the conversational context  Han et al. (2022); Chung et al. (2021). Table 1 shows an example of counterspeech generated by prompt with Llama2. The counterspeech is perceived to be robotic, as it ignores the hatred and anger directed at the victim in the language (“racist feminist c*** …”) and instead focuses on criticizing the imagined scenario (“not put it out if she were on fire”). The final suggestion is vague and does not specifically address the harmful stereotypes in the original statement. It may lack nuanced understanding and the depth of empathy for effective communication Hangartner et al. (2021).
@@
Human likeness is an important factor in crafting counterspeech. Studies find that more human-like counterspeech tends to be more effective Pinochet et al. (2024). The tone and form of counterspeech significantly affect the effectiveness in influencing behaviors of users Buerger (2021). Generated counterspeech that embodies human characteristics including empathy, humor, and respect can foster emotional attachments Glikson and Woolley (2020), potentially making counterspeech more acceptable and reliable Jiang et al. (2023b). Human-like responses are more likely to express understanding and social connection Go and Sundar (2019a), contribute to fostering genuine interactions, and build trust Fenwick and Molnar (2022).
@@
On the other hand, there are ethical concerns if AI-generated and human-written counterspeech are indistinguishable Hua et al. (2024). AI models may be used for manipulative purposes such as swaying opinions not aligned with specific agendas Karnouskos (2020). These models may also generate biased responses that could unintentionally reinforce stereotypes or cause harm Lucy and Bamman (2021); Kenthapadi et al. (2023). Evaluating the human likeness of AI-generated counterspeech is important because it brings awareness to these issues and could help mitigate them.
@@
Previous studies have focused on evaluating the relevance, quality, toxicity, persuasiveness, and effectiveness of counterspeech Saha et al. (2022); Hong et al. (2024). Some studied linguistic aspects using metrics such as BLEU, METEOR, BERTScore, GRUENV scores, type-token and distinct-n Zhu and Bhat (2021); Tekiroglu et al. (2022). However, few have conducted thorough evaluations regarding human likeness.
@@
We implement several state-of-the-art counterspeech generation strategies with LLMs, including prompting for one counterspeech (Prompt), prompt for multiple replies and select the best(Prompt and Select), fine-tune LLMs with existing counterspeech datasets(Fine-tune), and outcome-constrained LLMs.(Constrained). We also collect human-written counterspeech, including both examples from actual user-generated content (i.e., genuine content from Reddit) and crowd workers (i.e., counterspeech generated on demand).
With AI-generated and human-written counterspeech, we build authorship attribution models to identify which AI models generate more human-like counterspeech. The higher the classification accuracy, the less human-like the AI-generated counterspeech is. We further deploy a human annotation task to validate and check whether humans can discern the differences between AI-generated and human-written counterspeech. Additionally, we analyze linguistic characteristics, politeness, and specificity of the counterspeech generated by various methods to understand the differences amongst counterspeech depending on the source (types of AI-generated and human-written).
Our study shows both classifiers and humans can easily distinguish AI-generated and human-written counterspeech. AI-generated counterspeech shows (a) significant differences in linguistic characteristics, and (b) is more polite and less specific.
@@
Domain Size Context? Generation Model Training Methods
@@ 2 Related Work
Counterspeech Generation Several studies employ crowd workers or experts to curate counterspeech Chung et al. (2019); Qian et al. (2019). These human-written datasets have been widely utilized for training counterspeech models Tekiroglu et al. (2022); Saha et al. (2022); Gupta et al. (2023). Hybrid approaches that combine human annotations and generative models have been proposed, resulting in counterspeech corpora such as CONAN Chung et al. (2019) and multiCONAN Fanton et al. (2021). Advanced generative models have been developed for a more sophisticated generation of counterspeech, for example, the prompt and select method has been shown to generate more diverse and relevant counterspeech Zhu and Bhat (2021). Other approaches include constraining models for generating polite Saha et al. (2022), intent-based Gupta et al. (2023), or outcome-oriented counterspeech Hong et al. (2024). This study implements several representative models to generate counterspeech and conducts an extensive evaluation focusing on human likeness.
@@
Most curated datasets consist of texts without additional context El-Sayed and Nasr (2023); Buz et al. (2024); An et al. (2023). Few include context and involve sequences or groups of related texts Ji et al. (2024); Prajapati et al. (2024). To distinguish between AI-generated and human-written texts, many researchers conduct linguistic analysis and authorship attribution, supplemented with human evaluations Zhou et al. (2023); Ma et al. (2023); Buz et al. (2024); Ji et al. (2024); Prajapati et al. (2024). Our study is among the few to work in a dialogue setting, where texts are dialogue turns countering a hate speech—AI-generated following four strategies or human-written from two sources.
3 Counterspeech Data
@@ 3.1 Human-written Counterspeech
Human-written counterspeech consists of replies to hateful content written by humans. We consider counterspeech written by both genuine Reddit users (i.e., users that post out of their own will) and crowd workers who are tasked with writing counterspeech. Genuine human-written counterspeech is collected from Reddit. First, we use keyword- and community-based sampling methods to retrieve hate speech posts from 42 subreddits (Appendix A) with a higher prevalence of hate via the Pushshift API  [^2] https://pushshift.io/api-parameters/. This step results in 27,491 hate speech posts and their replies. Second, we identify replies that are counterspeech with three BERT classifiers individually fine-tuned with three existing counterspeech corpora Qian et al. (2019); Chung et al. (2021); Yu et al. (2022). We consider a reply to be counterspeech if the three classifiers indicate so, finally resulting in 14,973 (hate speech, counterspeech) pairs from Reddit.
Counterspeech written by crowd workers is collected from the Benchmark dataset Qian et al. (2019). This corpus includes hate speech posts paired with counterspeech replies written by crowd workers on demand. We identify 14,208 valid (hate speech, counterspeech) pairs in Benchmark.
@@ 4 Evaluation Methods
We conduct evaluations comparing human-written and AI-generated counterspeech accounting for four major categories: human likeness, linguistic differences, politeness, and specificity.
@@
If it is easy to differentiate between human-written and AI-generated counterspeech, we can conclude that the latter is not human-like. We develop BERT classifiers with our curated dataset (Section 3). The dataset is divided into an 80/20 split for training and testing. Specifically, we create five binary classifiers (human-written or AI-generated) depending on what is included in the AI-generated counterspeech: all or only the counterspeech generated by one of the four strategies (see Appendix B for details).
In addition, we conduct a human validation for deeper insights. We randomly select 100 samples from each AI strategy (4×100)4$\times$ 100)4 × 100 ) and human-written counterspeech (2×100)2$\times$ 100)2 × 100 ), combine them, and ask human annotators to label whether the counterspeech is human-written or AI-generated. Two research assistants complete the annotation process. Table 3 presents the agreement rate and Cohen’s Kappa score. Both indicate high reliability, meaning that the task is straightforward for humans.
@@
The level of politeness assesses the degree of respectfulness and courtesy. It provides a complementary perspective to understand the differences between AI-generated and human-written counterspeech. We build a politeness prediction model, a BERT model fine-tuned with the dataset by Saha et al. (2022). The model achieves an F1 score of 0.91. We use this model to predict the politeness level of the counterspeech replies on a scale of 0 to 7. A higher score indicates more politeness.
@@ 5.4 Specificity
The counterspeech generated by Prompt and Prompt and Select models exhibits a similar distribution with a moderate specificity level of 3 (Figure 2(b)). The Constrained model-generated counterspeech has a median specificity score of 1.5, indicating that it lacks specificity and provides a more general reply. Comparatively, the counterspeech by Fine-tune models shows high variances in the specificity scores, but most scores are between 3 and 4, indicating counterspeech is more targeted to the hate speech post at hand.
@@ 7 Conclusion
We fully consider the potential risks and benefits of our study. First, we collect data from Reddit, a public forum that makes data available to third parties. We have masked users’ names and identities before analysis. Second, our study employs human annotators to evaluate counterspeech, their names and identities are encrypted to avoid the identification of annotators. Third, we provide various AI-based methods to generate counterspeech that embody human likeness. There is a concern that some generated responses are indistinguishable from those created by humans, raising ethical considerations. We acknowledge the potential risks these methods may cause. However, the benefits will outweigh such risks, for example, using these generations in adversarial learning to develop more robust models for identifying AI-generated content.
@@ Acknowledgements
This work was supported by the Institute of Museum and Library Services (IMLS) National Leadership Grants under LG256661-OLS-24 and LG-256666-OLS-24.
@@ Appendix C Annoatation Details
0-Very Impolite Counterspeech is perceived as rude or unfriendly The study was b**.